OpenStudy (drewkatski):

Can someone please clarify this? I don't understand the steps to get the answer. If f (x) = 3x^2-4 Write f (2x) in its simplest form

OpenStudy (emmigrace222):

http://www.purplemath.com/modules/fcnops.htm :) try that

OpenStudy (mathmale):

Hello, Katski, You're given a function: f (x) = 3x^2 - 4, and you're asked to evaluate this function when the input is 2x instead of just x. You might want to re-write the given function as f( ) = 3( )^2 - 4 and then write "2x" inside the parentheses. Simplify the result, and you've then got your answer.

OpenStudy (mathmale):

thanks, D-L. My only suggestion is that you evaluate the function with 2x as input, not just 2 as input.

OpenStudy (drewkatski):

Hmm I see, so f (x) = 3 (2x)^2 - 4 ? Then i just work it out as normal? @mathmale

OpenStudy (drewkatski):

*f (2x) = ********

OpenStudy (mathmale):

f (2x) = 3 (2x)^2 - 4. Yes. You must do exponentiation BEFORE multiplication, so focus on (2x)^2. That equals what?

OpenStudy (drewkatski):

mmm, 4x^2

OpenStudy (mathmale):

Great. Now re-write f (2x) = 3 (2x)^2 - 4 using your latest result.

OpenStudy (drewkatski):

f (2x) = 3(4x^2) - 4 i think

OpenStudy (drewkatski):

@mathmale

OpenStudy (whpalmer4):

Yes, that is correct, but simplify/expand it all the way.

OpenStudy (mathmale):

Drew: Just expand your f (2x) = 3(4x^2) - 4: f(2x) = 12x^2-4. that's it.